# Fabien Le Floc’h

### v1

@@ 3.2 First order expansion
(18) $V_{\textmd{VG1}}=A_{p}\cdot\mathrm{Black}(1,\kappa^{\star},\nu^{2},T)\\ +A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\kappa^{\star}A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(1,\kappa^{\star},\nu^{2},T\right)\\ -A_{p}\sum_{i=1}^{n}\tilde{a}_{i}^{\star}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\,.$
@@ 3.3 Second order expansion
(19) $V_{\textmd{VG2}}=V_{\textmd{VG1}}+\frac{A_{p}e^{\nu^{2}(T)}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)+\frac{\kappa^{2}A_{p}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\kappa^{\star}A_{p}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\kappa^{\star}{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$
with $\bar{v}_{i}^{-}$, ${v}_{i,j}$ defined in Equations 36 and 31.
@@ 3.4 Third order expansion
(20) $V_{\textmd{VG3}}=V_{\textmd{VG2}}\\ +\frac{A_{p}e^{3\nu^{2}(T)}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{3\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}\kappa^{\star}e^{\nu^{2}}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}{\kappa^{\star}}^{2}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}{\kappa^{\star}}^{3}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\frac{A_{p}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu^{2}+2\nu_{n}^{2}-2\nu_{pn}^{2}+2\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}{\kappa^{\star}}^{2}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +A_{p}{\kappa^{\star}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{3\nu_{n}^{2}-2\nu_{pn}^{2}+\bar{v}_{i}+\bar{v}_{j}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\nu^{2}-2\nu_{pn}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}\kappa^{\star}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}}{6}\sum_{i,j,k=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\tilde{a}_{k}^{\star}e^{{v}_{i,j}+{v}_{k,i}+{v}_{k,j}+2\bar{v}_{i}^{-}+2\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+3\nu_{n}^{2}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{k}-3\nu_{np}^{2}+3\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$

### v2

@@ 3.2 First order expansion
(18) $V_{\textsf{VG1}}=A_{p}\cdot\mathrm{Black}(1,\kappa^{\star},\nu^{2},T)\\ +A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\kappa^{\star}A_{p}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(1,\kappa^{\star},\nu^{2},T\right)\\ -A_{p}\sum_{i=1}^{n}\tilde{a}_{i}^{\star}\frac{\partial}{\partial\kappa^{\star}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}(T)-\nu_{n}^{2}(T)-\nu^{2}(T)}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\,.$
@@ 3.3 Second order expansion
(19) $V_{\textsf{VG2}}=V_{\textsf{VG1}}+\frac{A_{p}e^{\nu^{2}(T)}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)+\frac{\kappa^{2}A_{p}}{2}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\kappa^{\star}A_{p}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\kappa^{\star}{A_{p}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{2}}{\partial{K^{\star}}^{2}}\mathrm{Black}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{n}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{2}}{\partial{\kappa^{\star}}^{2}}\mathrm{Black}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$
with $\bar{v}_{i}^{-}$, ${v}_{i,j}$ defined in Equations 36 and LABEL:eqn:vij.
@@ 3.4 Third order expansion
(20) $V_{\textsf{VG3}}=V_{\textsf{VG2}}\\ +\frac{A_{p}e^{3\nu^{2}(T)}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{3\nu^{2}(T)},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}\kappa^{\star}e^{\nu^{2}}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}{\kappa^{\star}}^{2}}{2}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}},\kappa^{\star},\nu^{2},T\right)-\frac{A_{p}{\kappa^{\star}}^{3}}{6}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(1,\kappa^{\star},{\nu}^{2},T\right)\\ -\frac{A_{p}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu^{2}+2\nu_{n}^{2}-2\nu_{pn}^{2}+2\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{2\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}{\kappa^{\star}}^{2}}{2}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{-\frac{\nu_{p}^{2}-\nu_{n}^{2}-\nu^{2}}{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +A_{p}{\kappa^{\star}}\sum_{i=1}^{m}\tilde{a}_{i}^{\star}e^{\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\nu^{2}+\nu_{n}^{2}-\nu_{pn}^{2}+\bar{v}_{i}},\kappa^{\star},\nu^{2},T\right)\\ +\frac{A_{p}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{3\nu_{n}^{2}-2\nu_{pn}^{2}+\bar{v}_{i}+\bar{v}_{j}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\nu^{2}-2\nu_{pn}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}\kappa^{\star}}{2}\sum_{i,j=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}e^{\nu_{n}^{2}+{v}_{i,j}+\bar{v}_{i}^{-}+\bar{v}_{j}^{-}(T)}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}-2\nu_{np}^{2}+2\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\\ -\frac{A_{p}}{6}\sum_{i,j,k=1}^{m}\tilde{a}_{i}^{\star}\tilde{a}_{j}^{\star}\tilde{a}_{k}^{\star}e^{{v}_{i,j}+{v}_{k,i}+{v}_{k,j}+2\bar{v}_{i}^{-}+2\bar{v}_{j}^{-}+2\bar{v}_{k}^{-}+3\nu_{n}^{2}}\frac{\partial^{3}\mathrm{Black}}{\partial{\kappa^{\star}}^{3}}\left(e^{\bar{v}_{i}+\bar{v}_{j}+\bar{v}_{k}-3\nu_{np}^{2}+3\nu_{n}^{2}},\kappa^{\star},\nu^{2},T\right)\,,$